UT Jan 15 2022 - Split Night LBTI (1st half) / DEC (2nd half)



Night Info

  • Sunset 17:41
  • Twilight Ends 19:02
  • Night Center 00:29
  • Moonset 05:50
  • Twilight Begins 05:55
  • Sunrise 07:17

Night Plan

  • LBTI until Night Center - 00:29
  • LBTI Eng Time 2hr - 00:29-02:29
    • Obtain photometric standards in most LMIRCam filters
    • Seeing up to 2"
    • Thin clouds OK
  • Reconfigure from LBTI to PEPSI-Pol 02:30-02:42
  • PEPSIPOL Eng - 1hr 02:42-03:42
  • DONE PEPSIPOL Eng backup if clouds are thicker or seeing > 1.2
    • use IDL zernike_collect for trefoil and quadrafoil (not sure if present setup will handle quadrafoil)
    • can be done on a substantially brighter star if there are clouds
    • use z9_10.list for trefoil, use z14_15.list for quadrafoil - beware z14 and z15 might be turned off.
  • IF LFBG is functional (It is NOT functional.)
    • Reconfigure from PEPSI-Pol to LUCI-Bino 03:42-03:54
    • LUCI-1 AO/ESM checkouts 03:54-twilight
      • Bin 3 and Bin 4 Full AO
      • ESM Bin 2 and Bin 4 - take data with SL LUCI-2 in parallel
    • Seeing < 1.2" for full AO, Seeing < 2" for ESM
    • Thin Clouds OK
  • IF LFBG is still offline
    • Reconfigure from PEPSI-Pol to LUCI-Bino 03:42-03:54 - unless J.Hill needs additional time for PEPSI-Pol eng work. This PEPSI POL work has higher priority
    • Imaging of GRB211211A - LUCI-2
      • /home/lbto/Service_Observing/2021B_D_E/LUCI/GRB211211A/Grb211211A_LD2021B-002_OT.xml
      • Ks imaging - 1.25hr w/overheads
      • Seeing < 1.2" and thin clouds ok
  • Poor Weather Backup
    • For LBTI w/seeing >2" and/or thick clouds and/or PEPSI-Pol time and/or LUCI-1 AO can't be executed
      • Reconfigure to LUCI-2
      • Local Mergers - Rothberg
        • /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers
        • OT file: COBandhead_Mergers_V37.xml
        • No specific priorities - pick whats visible and avoids wind, clouds, etc...
          • Arp 156 midnight-twilight - 1hr (more if super thick clouds or super poor seeing)
          • NGC 3597 - 02:30-05:30 - 1hr (more if super thick clouds or super poor seeing)
          • UGC 9829 - 04:30-twilight - 1hr (more if super thick clouds or super poor seeing)
          • UGC 10607 (use earlier telluric) - 05:00-twilight - 1hr
          • UGC 10675 (use earlier telluric - same as UGC 10607) - 05:00-twlight - 1hr

Overview

LBTI science first half. Second half mostly lost to weather. Opened for 1.5 hours at the end to get some zernike_collect in poor seeing with PEPSIPOL.

Staff

  • OSA: D. Huerta
  • ISA: J. Power

Night Details

23:35 Afternoon checkout of LUCI and other. David is configuring for LBTI SX and executing test preset for LBTI science is first half. 0.5" longslit put in LUCI2 and field stop alignment verified.

LBTI Science

LBTI science first half (see LBTI logs)



7:00 UT David has closed for high winds. Seeing prior was 2.5+". LBTI has said that even if we reopen they will not execute photometric standards with this poor seeing.

7:14 LBTI has handed over for engineering. We are closed for high winds. Reconfigured for PEPSI POL. Will execute the zernike collection if we open.

9:20 Winds appear to be trending downward

9:44 Winds back up gusting 22-22.5m/s

11:54 Winds are holding right below 20m/s so we are going to give it a try.

zernike_collect of Z9/Z10 for PEPSIPOL in poor seeing

12:00 UT Opening and Slewing for pointing and collimation. Initial pointing and collimation was ugly. Very out of focus and DX was way off. Need to go to much brighter star.

Sat Jan 15 12:08:52.072 2022 left Unauthorized BinocularControl command rejected for MODS at directGregorian focus
But unknown why David got that error.

12:18 Pointed and collimated on 13.8 magnitude pointing star at RA 12:15:45.4, 14:27:29.93. Seeing 1.45-1.55" on the left and 1.3-1.45" on the right.

presently IDL is set for LPOL/RPOL xy_transpose=0 (you don't need to know that - just for reference when processing data)

12:21 Starting zernike_collect with z9_10.list

Acquiring LEFT WFSC image (exp=15.0000 s)...
===> Error in gcs_take_image
err = Error: request for a full WFSCam image while WFS loop closed!
err = Command Failed.
David is resending preset in guide mode only. Light cirrus. Seeing 1.3" on the right and 1.6". We left image roughly collimated since if we did not it was severely out of focus. WFS off, but guiding.

zernike_collect on right fails:
===> Error in gcs_take_image
err = request to AzCamServer for full WFSCam failed. err=-1
err = Command Failed.

Restart DX azcam server, restart GCS, resend preset

12:38 David resent preset to both sided.

DX AZCam not up, DX AZCAm needed another kick. Restart successful this time.

12:47 Preset sent. David sent Active, so turned off WFS and have guiding only. Seeing is highly variable and poor 1.3-2" on both sides. For real starting the zernike_collect with z9_10.list

left_wfs000012-20

right_wfs000002-09

12:55 David is checking if z14 & 15 are off in PSF (Active Optics Gains). There was some question about those and whether we were doing this correctly. Calling John for clarification.

JMH (later): Yes, Z14/Z15 were turned off in IDL, so no on-the-fly analysis, but Z14/Z15 were applied to M1, and wfsc images were collected. Z9/Z10 data looks to be useable. Maybe we can recover Z14/Z15 in post-processing.

13:04 Executing zernike_collect with z14_15.list

left_wfs000021-29

right_wfs000010-18

grep "Primary Zernikes" 20220115.log | grep T13:0 | grep -v crosstalk

2022-01-15T13:04:04.097116+00:00 tcs2 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -1303.39 -2605.21 0 0 7.83408 70.8922 270.986 0 0 0 0 0 0 0 0 0 0 36.2851
2022-01-15T13:04:18.743275+00:00 tcs2 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -1304.1 -2609.49 0 0 7.51169 70.8922 270.986 0 0 200 0 0 0 0 0 0 0 36.2851
2022-01-15T13:04:21.825305+00:00 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 0 -217.134 -385.001 0 0 -70.6928 -56.7583 140.269 0 0 200 0 0 0 0 0 0 0 35.8928
2022-01-15T13:05:40.877716+00:00 tcs2 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -1308.14 -2633.67 0 0 5.69699 70.8922 270.986 0 0 0 0 0 0 0 0 0 0 36.2851
2022-01-15T13:05:48.669348+00:00 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 0 -217.134 -385.001 0 0 -70.6928 -56.7583 140.269 0 0 0 0 0 0 0 0 0 0 35.8928
2022-01-15T13:06:13.462369+00:00 tcs2 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -1309.74 -2643.33 0 0 4.97586 70.8922 270.986 0 0 0 200 0 0 0 0 0 0 36.2851
2022-01-15T13:06:20.375816+00:00 tcs1 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 0 -217.134 -385.001 0 0 -70.6928 -56.7583 140.269 0 0 0 200 0 0 0 0 0 0 35.8928
2022-01-15T13:07:37.878003+00:00 tcs2 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -1313.95 -2668.81 0 0 3.08286 70.8922 270.986 0 0 0 0 0 0 0 0 0 0 36.2851

13:16 slewing to another target. Sunrise is approaching so we will rerun this on a secondary target and monitor logs. Mag 11.7 (seeing 2" and some cirrus). Preset sent in guide mode

zernike_collect with z9_z10.list

left_wfs000030-38

right_wfs000019-27

13:23 starting zernike_collect with z14_z15.list. Seeing 2-2.9" and Cirrus.

left_wfs000039-47

right_wfs000028-36

13:28 Seeing was so bad, but seems to have settled down a bit. Will rerun quadrafoil. Seeing ~1.4"

Executing zernike_collect with z14_15.list

left_wfs000048-56

right_wfs000037-45

13:36 Mag 10.6 target, el 66 (RA12:11:11, dec 33:15:39) guide preset. Cirrus,seeing improving a bit. Sun is coming up but sneaking one more in. Seeing 1.2-1.6".

zernike_collect with z9_10.list

left_wfs000057-65

right_wfs000046-54

zernike_collect with z14_15.list

left_wfs000066-74

right_wfs000055-63


zernike_collect analysis

I used IDL wfsc_process_image.pro to reprocess the zernike_collect data from 20220115 with a variety to values of the Zernike rotation offset (wfsc.rot_offset). This is a simple rotation which could be done in a spreadsheet, but we use IDL preprocessing to make sure the signs are consistent. Because the same data set of 3-4 measurements of 3 images per Zernike term is used, the rotations make a perfect sine wave. Any noise from seeing variations shows up as differences between the individual plots in terms of the preferred angle. The rms between sets of 3 measurements is about 20% of the peak values (40 nm rms for a mean of 200 nm).

This discussion will include results from the SX side (LPOL@AGW3), but the DX side is similar with an offset of 1.5 degrees.

The first change required was to change the signs of Z14/Z15 relative to each other. The attached plots have a -1 applied to Z15.

Next, we notice that Z9/Z10 and Z14/Z15 don't prefer the same rotation offset angle. Z9 likes 129, Z10 likes 135 for an average of 132. Z14 likes 117, Z10 likes 118 for an average of 117.5. We assume this is because trefoil and quadrafoil are on different lobes of their pattern of degenerate rotation.

If we gave quadrafoil Z14/Z15 a 90 degree degenerate rotation, the angle would be 118-90=28 or 118+90=208 which could extend to 298. If we changed the signs of both Z14/Z15 to -1/+1, then the good angle would move to either 118+45=163 or 118-45=73 which could extend to 253. With degenerate rotations and sign flips, Quadrafoil can be 28,73,118,163,208,253,298,343 deg.

If we gave trefoil Z9/Z10 a 120 degree degenerate rotation, the angle would be 132-120 = 12 or 132+120=252. Trefoil can be 12,72,132,182,252,312 deg.

Can the trefoil/quadrafoil common ground of 72 or 252 be consistent with astigmatism and coma? I don't see that right now.....

We know that ~146 is at least semi-stable for both astigmatism and coma. Going back to read the 20211227 UT wiki again we see that coma and astigmatism seem stable-ish between 124 and 146 degrees giving the three sets of possible solutions below.

That means changing sign of both astigmatisms would put us at 146+90=236 (within 22 deg of 252) or at 146-90=56, or a 180 deg rotation at 146+180=326.
Astigmatism can be 56,146,236,326 with a +-22 deg range for stability.
Astigmatism can be 40,132,222,312 with a +-22 deg range for stability.
Astigmatism can be 32,124,214,304 with a +-22 deg range for stability.

That means changing sign of both comas would put us at 146+180=326 (not within 45 deg of 252). Thus it seems coma is NOT in the range of compatible angles.
Coma can be 146,326 with +-45 degree range for stability.
Coma can be 132,312 with +-45 degree range for stability.
Coma can be 124,318 with +-45 degree range for stability.

JMH is presently perplexed about this angular dilemma. Of course, if we don't ever use quadrafoil, the angles may all be OK (but not optimized) near 132.

Change by Coma Astig Trefoil Quadrafoil
Degenerate 360 180 120 90
Opposite 180 90 60 45
Crossed 90 45 30 22.5
Spiral-in 45 22.5 15 11.25

  • LPOL 400 nm Z9 applied:
    LPOL 400 nm Z9 applied.png

  • LPOL 400 nm Z10 applied:
    LPOL 400 nm Z10 applied.png

  • LPOL 400 nm Z14 applied:
    LPOL 200 nm Z14 applied.png

  • LPOL 400 nm Z15 applied:
    LPOL 200 nm Z15 applied.png

  • RPOL 400 nm Z9 applied:
    RPOL 400 nm Z9 applied.png

  • RPOL 400 nm Z10 applied:
    RPOL 400 nm Z10 applied.png

  • RPOL 400 nm Z14 applied:
    RPOL 200 nm Z14 applied.png

  • RPOL 400 nm Z15 applied:
    RPOL 200 nm Z15 applied.png

-- BarryRothberg - 11 Jan 2022
I Attachment Action Size Date Who Comment
LPOL 200 nm Z14 applied.pngpng LPOL 200 nm Z14 applied.png manage 21 K 24 Jan 2022 - 21:51 JohnHill LPOL 400 nm Z14 applied
LPOL 200 nm Z15 applied.pngpng LPOL 200 nm Z15 applied.png manage 21 K 24 Jan 2022 - 21:52 JohnHill LPOL 400 nm Z15 applied
LPOL 400 nm Z10 applied.pngpng LPOL 400 nm Z10 applied.png manage 20 K 24 Jan 2022 - 21:50 JohnHill LPOL 400 nm Z10 applied
LPOL 400 nm Z9 applied.pngpng LPOL 400 nm Z9 applied.png manage 20 K 24 Jan 2022 - 21:49 JohnHill LPOL 400 nm Z9 applied
RPOL 200 nm Z14 applied.pngpng RPOL 200 nm Z14 applied.png manage 21 K 24 Jan 2022 - 21:56 JohnHill RPOL 400 nm Z14 applied
RPOL 200 nm Z15 applied.pngpng RPOL 200 nm Z15 applied.png manage 21 K 24 Jan 2022 - 21:56 JohnHill RPOL 400 nm Z15 applied
RPOL 400 nm Z10 applied.pngpng RPOL 400 nm Z10 applied.png manage 20 K 24 Jan 2022 - 21:53 JohnHill RPOL 400 nm Z10 applied
RPOL 400 nm Z9 applied.pngpng RPOL 400 nm Z9 applied.png manage 19 K 24 Jan 2022 - 21:53 JohnHill RPOL 400 nm Z9 applied
Topic revision: r12 - 24 Jan 2022, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback